### v1

@@ Abstract
Keywords: Score-driven models, Rough volatility, Weak convergence, Option pricing
2020 MSC: 60F17, 60G22, 62M10, 91B70
@@ 1 Introduction
Score-driven models, also known as generalized autoregressive score models, were proposed by Creal et al., (2013) and Harvey, (2013) as a class of observation-driven time series models in which parameter updates are driven by the score—the gradient of the log-likelihood function. By directly linking the score of the data distribution to parameter dynamics in a unified framework, these models inherit flexibility and generality, encompassing many classical models such as GARCH (Bollerslev, , 1986) and ACD (Engle and Russell, , 1998) models. Theoretically, score-driven models have been proven optimal in minimizing the Kullback-Leibler divergence between the true distribution and the postulated distribution (Gorgi et al., , 2024). Moreover, in statistical inference, the model’s closed-form likelihood function facilitates efficient maximum likelihood estimation, exhibiting lower computational costs compared to parameter-driven models. In application, they also demonstrate robust empirical performance across diverse fields such as economics, finance and biology. Within just a few years, the literature on SD models has grown to nearly 400 and can be found online at http://www.gasmodel.com/gaspapers.htm.
@@
As a remarkably influential model in time series analysis, it naturally raises the question: what would its continuous-time counterpart look like? Buccheri et al., (2021) have investigated its continuous-time limit, obtaining a two-dimensional diffusion process. Similarly, Wu and He, (2024) investigated the continuous-time limit of the quasi score-driven volatility model proposed by Blasques et al., (2023), and derived a correlated stochastic volatility model. However, the score-driven models they studied only incorporate first-order or finite-order lags—these are all Markov processes. In fact, when the model was first introduced in Creal et al., (2013), the question was raised as to whether it could be extended to an infinite-lag form. Janus et al., (2014) and Opschoor and Lucas, (2019) pursued such extensions by proposing long memory score-driven models. This naturally leads to the question: What would be the continuous-time counterpart of such a long memory version?
@@
On the other hand, since Gatheral et al., (2018) discovered that volatility is rough, rough volatility models have gained popularity in financial modeling. These models are characterized by the use of fractional Brownian motion with a Hurst parameter $H<1/2$, whose paths are rougher than those of Brownian motion—hence the name. El Euch et al., (2018) investigated the microstructure of rough volatility by employing Hawkes processes to model the arrival dynamics in limit order books. This theoretical basis originates from the discovery of Jaisson and Rosenbaum, (2016) that the scaling limit of nearly unstable heavy-tailed Hawkes processes is a rough Cox-Ingersoll-Ross (CIR) process. It is worth noting that Hawkes processes are self-exciting (similar to observation-driven mechanisms) and exhibit long memory, which provides inspiration for us. In particular, Cai et al., (2024) studied the discrete-time counterpart of the Hawkes process—the INAR($\infty$) process—and obtained similar convergence results. This motivates our investigation into the scaling limit of score-driven models when extended to incorporate long memory.
@@
This paper establishes that a sequence of long memory score-driven models, under appropriate scaling, converges weakly to a rough Ornstein-Uhlenbeck (OU) process. By modeling time-varying parameters as log-volatility (as in EGARCH-type models), we obtain the limit a class of rough volatility models. Consequently, long memory score-driven volatility models can serve as discrete-time approximations for such rough volatility models, facilitating efficient Monte Carlo simulations for financial applications. To the best of our knowledge, apart from Wang and Cui, (2025)’s approach using INAR($\infty$) processes to approximate the rough Heston model, this study presents one of the few methodologies employing time series models to approximate rough volatility processes. Moreover, we present specific examples of long memory score-driven volatility models to further demonstrate the practical versatility and broad applicability of our theoretical results.
@@
The remainder of the paper is structured as follows. Section 2 introduces the long memory score-driven models that we focus on and provides a heuristic derivation of their continuous-time limits. In Section 3, we establish the main convergence results and presents their proofs. Section 4 applies these results to volatility modeling, illustrating with the Gamma-GED-EGARCH($\infty$) model and its extension. We conducts corresponding approximations of rough volatility models and option pricing. Section 5 concludes our findings and provides an outlook for future research.
@@ 2 Long memory score-driven models
Since Engle, (1982) discovered that economic data exhibit conditional heteroskedasticity, it has gradually been recognized that certain characteristics of time series are not immutable. Specifically, let $\{y_{n}\}_{n\in\mathbb{N}}$ be a time series and $\mathcal{F}{n}=\sigma(y_{n},y_{n-1},\ldots,y_{0})$ be the $\sigma$-algebra it generates. The conditional density
@@
with a time-varying parameter $\lambda_{n}$. To address this issue, Cox et al., (1981) proposed two modeling approaches: observation-driven models, where the dynamics of the time-varying parameters are governed by past observations of the series itself, and parameter-driven models, in which the parameter dynamics are driven by another stochastic process.
@@
$\lambda_{n+1}=\omega+\beta\lambda_{n}+\alpha S(\lambda_{n})\nabla_{n}.$ (2.1)
where $\nabla_{n}=\frac{\partial\log p(y_{n}|\lambda_{n},\mathcal{F}{n-1})}{\partial\lambda_{n}}$ denotes the partial derivative of the log-density function of observations with respect to the parameter, commonly referred to as the (Fisher) score in statistics. Consequently, these models are aptly termed score-driven models. The function $S(\cdot)$ is known as the scaling function, which accounts for the curvature of the log-density function. It is typically chosen as the negative exponent of the conditional Fisher information, i.e., $S(\lambda_{n})=[\mathbb{E}(\nabla_{n}^{2}|\mathcal{F}{n-1})]^{-a}$, with common choices for $a$ being 0, $1/2$, or $1$. To illustrate, when the density function $p(\cdot)$ is chosen to be that of a normal distribution and the parameter $\lambda$ is identified as the conditional variance $\sigma^{2}$, with $a=1$, the score-driven model reduces to the classical GARCH model.
The equation given by (2.1) represents a first-order score-driven model, as the relationship between the score and the parameter is lagged by only one period. This framework can be extended to higher orders, such as $p$-th order or even infinite order, as in Janus et al., (2014),
@@
Noting that our time index starts from 0, the summation at time $n$ includes only $n$ terms. The term “infinite” refers to the dependence on the entire history. Additionally, we omit the autoregressive term, as it can be absorbed into the coefficients $\phi_{i}$ through the moving average representation. This formulation appears similar to the ARCH($\infty$) model, but the key difference lies in the fact that the infinite-order lagged term $S\nabla$ in the score-driven model is a martingale difference, whereas the $y^{2}$ term in the ARCH($\infty$) model is not. In other words, rewriting the ARCH($\infty$) model in a form driven by martingale differences gives
@@
which includes an additional infinite autoregressive component. To address this, Opschoor and Lucas, (2019) proposed the following formulation:
@@
where $\phi_{i}^{(n)}=a_{n}\phi_{i}>0$. Let $a_{n}>0$ and $\uparrow 1$ as $n\to\infty$, $\|\phi\|=\sum_{i=1}^{\infty}\phi_{i}=1$, so that $\|\phi^{(n)}\|\uparrow 1$. This case is referred to as nearly unstable. Specifically, we set $\phi_{i}\sim K/i^{1+\alpha}$, exhibiting power-law decay, where $\alpha\in(1/2,1)$, a form referred to as “heavy-tailed” in Jaisson and Rosenbaum, (2016). This condition is critical for the limit to generate roughness. The sigma-algebra is defined as $\mathcal{F}^{(n)}_{t}=\sigma(\lambda_{0}^{(n)},y_{i}^{(n)},\lambda_{i+1}^{(n)},i=0,1,\ldots,t)$, since $\lambda_{t+1}^{(n)}$ is determined when $\{y_{i}^{(n)}\}_{i\leq t}$ are known.
@@
where $\{W_{t}\}_{t\geq 0}$ is a standard Brownian motion. And according to Cai et al., (2024), in that case $1/2<\alpha<1$, when $1-a_{n}\sim n^{-\alpha}$ tends to zero and the coefficients $\phi_{i}\sim K/i^{1+\alpha}$ exhibit power-law decay, the following weak convergence holds:
@@
In fact, similar to Proposition 4.10 in El Euch et al., (2018), the SDE (2.7) has the same solution as the SDE
@@
Since $\alpha-1\in(-1/2,0)$, the term $\frac{1}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}{\rm d}W_{s}$ corresponds to a fractional Brownian motion with Hurst parameter $H<1/2$, which means (2.8) falls into the rough case. It is worth mentioning that when $U(x)=\sqrt{x}$, this actually corresponds to the rough fractional CIR process described in Jaisson and Rosenbaum, (2016).
@@ 3 Main results
There exists a constant $\gamma<\infty$, such that $U(x)\to\gamma$, as $x\to\infty$.
@@
For each $n$, the martingale difference sequence $\{\Delta M_{k}^{(n)}\}_{k\in\mathbb{N}}$ is strictly stationary. Furthermore, for any $p\geq 2$,
$\sup_{n\in\mathbb{N}}\mathbb{E}\left|\Delta M_{k}^{(n)}\right|^{p}<\infty.$
@@
In fact, we restrict our attention to the case of rough OU process in (2.8). This is because our proof strategy fundamentally differs from that of Jaisson and Rosenbaum, (2016) or Wang and Cui, (2025): while they show that the Hawkes process (equivalent to the integrated intensity in probability) converges to a integrated rough CIR process, we directly establish the convergence of the parameter process itself to a rough process, rather than its integral. The difficulty in dealing with the process itself lies in verifying its tightness and we attempt to circumvent this by applying the continuous mapping theorem.
To this end, we regard the rough diffusion process as the image of a diffusion process under the generalized fractional operator (GFO), which is inspired by Horvath et al., (2024). Specifically, for $\lambda\in(0,1)$ and $\alpha\in(-\lambda,1-\lambda)$, $\mathcal{G}^{\alpha}$ is the GFO associated to kernel
@@
$(\mathcal{G}^{\alpha}f)(t)=\left\{\begin{aligned} &\int_{0}^{t}(f(s)-f(0))\frac{{\rm d}}{{\rm d}t}g(t-s){\rm d}s,\quad\text{if}\ \alpha\geq 0,\\ &\frac{{\rm d}}{{\rm d}t}\int_{0}^{t}(f(s)-f(0))g(t-s){\rm d}s,\quad\text{if}\ \alpha<0.\end{aligned}\right$.
When $g(x)=x^{\alpha}$, this reduces to the classical Riemann-Liouville operator. In this paper, we choose $g(x)=f^{\alpha,\kappa}(x)\in\mathcal{L}^{\alpha-1}$ and focus on the case $\alpha\in(1/2,1)$. Consequently, the corresponding GFO actually becomes
@@
$\Lambda_{t}=\theta\int_{0}^{t}f^{\alpha,\kappa}(s){\rm d}s+\nu\gamma\int_{0}^{t}f^{\alpha,\kappa}(t-s){\rm d}W_{s}$
is actually has the same solution as the (3.1). Consider the partition $\mathcal{T}:=\{t_{i}=i/n,i=0,1,\ldots,n\}$. We then define
@@ 3.1 Converges to $f^{\alpha,\kappa}(x)$
Jaisson and Rosenbaum, (2016) prove that the mean of geometric sums converges in distribution to a random variable with density function $f^{\alpha,\kappa}(x)$. Our setting differs slightly because the summed terms follow a discrete distribution.
@@
where $\hat{\phi}$ denotes the Fourier transform of $\phi$, or equivalently, the characteristic function of the random variable $U_{1}$. Since $\phi_{i}\sim\frac{K}{i^{1+\alpha}}$, by the Karamata-Tauberian theorem (see Cai et al., (2024)), we can characterize the behavior of $\hat{\phi}$ near the origin,
@@
Since $\lim_{x\to\infty}\phi_{n}(x)=0$, integration by parts yields
Since $\|\phi\|=1$, for any $|u|>1$, we have
$|1-\hat{\phi}_{n}(u)|\geq m\left|\frac{u}{n}\right|^{\alpha}.$
$|1-\hat{\phi}_{n}(u)|\geq m^{\prime}.$
Consequently, for $|u|>1$, taking $c_{2}=\min\left\{1/2,m^{\prime}\right\}$ yields the desired result. ∎
@@
$\|\rho_{n}-f^{\alpha,\kappa}\|_{2}:=\int_{0}^{1}\left[\rho_{n}(x)-f^{\alpha,\kappa}(x)\right]^{2}{\rm d}x\to 0,\ \text{as}\ n\to\infty.$ (3.7)
@@
To establish the $L^{2}$-convergence, we employ the Fourier isometry
According to estimates in Lemma 2, when $|u|\leq 1$,
when $|u|>1$,
$|\hat{\rho}_{n}(u)|\leq 1\wedge\frac{C}{|u|^{\alpha}},$
and the right-hand side
is $L^{2}$-integrable. The result thus follows by applying the dominated convergence theorem. ∎

### v2

@@ Abstract
keywords:
Score-driven models , Rough volatility , Weak convergence , Option pricing
2020 MSC:
60F17 , 60G22 , 62M10 , 91B70
journal: Journal of Econometrics
[…]
[label1]organization=School of Mathematics, Shanghai University of Finance and Economics,city=Shanghai, postcode=200433, country=China
@@ 1 Introduction
Score-driven models, also known as generalized autoregressive score models, were proposed by creal2013generalized and harvey2013dynamic as a class of observation-driven time series models in which parameter updates are driven by the score-the gradient of the log-likelihood function. By directly linking the score of the data distribution to parameter dynamics in a unified framework, these models inherit flexibility and generality, encompassing many classical models such as GARCH (bollerslev1986generalized) and ACD (engle1998autoregressive) models. Theoretically, score-driven models have been proven optimal in minimizing the Kullback-Leibler divergence between the true distribution and the postulated distribution (gorgi2023optimality). Moreover, in statistical inference, the model’s closed-form likelihood function facilitates efficient maximum likelihood estimation, exhibiting lower computational costs compared to parameter-driven models. In application, they also demonstrate robust empirical performance across diverse fields such as economics, finance and biology. Within just a few years, the literature on SD models has grown to nearly 400 and can be found online at gasmodel.com.
@@
As a remarkably influential model in time series analysis, one naturally raises the question: what would its continuous-time counterpart look like? Building on the method of nelson1990arch for deriving the continuous-time limit of GARCH, buccheri2021continuous have investigated continuous-time limit of score-driven volatility models, obtaining a two-dimensional diffusion process. Similarly, wu2024continuous investigated the continuous-time limit of the quasi score-driven volatility models proposed by blasques2023quasi, and derived a correlated stochastic volatility model. However, the score-driven models they studied only incorporate first-order or finite-order lags, which are all Markovian. In fact, when the model was first introduced in creal2013generalized, the question was raised as to whether it could be extended to an infinite-lag form. janus2014long and opschoor2019fractional pursued such extensions by proposing long memory score-driven models. This naturally leads to the question: What would be the continuous-time counterpart of such a long memory version?
@@
On the other hand, since gatheral2018volatility discovered that volatility is rough, rough volatility models have gained popularity in financial modeling. These models are characterized by the use of fractional Brownian motion with a Hurst parameter $H<1/2$, whose paths are rougher than those of Brownian motion, hence the name. el2018microstructural investigated the microstructure of rough volatility by employing Hawkes processes to model the arrival dynamics in limit order books. This theoretical basis originates from the discovery of jaisson2016rough that the scaling limit of nearly unstable heavy-tailed Hawkes processes is a rough Cox-Ingersoll-Ross (CIR) process. It is worth noting that Hawkes processes are self-exciting (similar to observation-driven mechanisms) and exhibit long memory, which provides inspiration for us. In particular, cai2024scaling studied the INAR($\infty$) process, a discrete-time counterpart of the Hawkes process, and obtained similar convergence results. This motivates our investigation into the scaling limit of score-driven models when extended to incorporate long memory.
@@
This paper establishes that a sequence of long memory score-driven models, under appropriate scaling, converges weakly to a rough Ornstein-Uhlenbeck (OU) process. By modeling time-varying parameters as log-volatility (as in EGARCH-type models), we obtain the limit a class of rough volatility models. Consequently, long memory score-driven volatility models can serve as discrete-time approximations for such rough volatility models, facilitating efficient Monte Carlo simulations for financial applications. To the best of our knowledge, apart from wang2025rough’s approach using INAR($\infty$) processes to approximate the rough Heston model, this study presents one of the few methodologies employing time series models to approximate rough volatility processes. Moreover, we present specific examples of long memory score-driven volatility models to further demonstrate the practical versatility and broad applicability of our theoretical results.
@@
The remainder of the paper is structured as follows. Section 2 introduces the long memory score-driven models that we focus on and provides a heuristic derivation of their continuous-time limits. In Section 3, we present our main convergence results. Proofs of the theorems are given in the main text when brief; all remaining proofs are deferred to the A. Section 4 applies these results to volatility modeling, illustrating with the Gamma-GED-EGARCH($\infty$) model and its extension, with the related proofs provided in B. We conducts corresponding approximations of rough volatility models and option pricing. Section 5 concludes our findings and provides an outlook for future research.
@@ 2 Long memory score-driven models
Since engle1982autoregressive discovered that economic data exhibit conditional heteroskedasticity, it has gradually been recognized that certain characteristics of time series are not immutable. Specifically, let $\{y_{n}\}_{n\in\mathbb{N}}$ be a time series and $\mathcal{F}_{n}=\sigma(y_{n},y_{n-1},\ldots,y_{0})$ be the $\sigma$-algebra it generates. The conditional density
@@
with a time-varying parameter $\lambda_{n}$. To address this issue, cox1981statistical proposed two modeling approaches: observation-driven models, where the dynamics of the time-varying parameters are governed by past observations of the series itself, and parameter-driven models, in which the parameter dynamics are driven by another stochastic process.
@@
$\lambda_{n+1}=\omega+\beta\lambda_{n}+\alpha S(\lambda_{n})\nabla_{n},$ (2.1)
where $\nabla_{n}=\frac{\partial\log p(y_{n}|\lambda_{n},\mathcal{F}{n-1})}{\partial\lambda_{n}}$ denotes the partial derivative of the log-density function of observations with respect to the parameter, commonly referred to as the (Fisher) score in statistics. Consequently, these models are aptly termed score-driven models. The function $S(\cdot)$ is known as the scaling function, which accounts for the curvature of the log-density function. It is typically chosen as the negative exponent of the conditional Fisher information, i.e., $S(\lambda_{n})=[\mathbb{E}(\nabla_{n}^{2}|\mathcal{F}_{n-1})]^{-a}$, with common choices for $a$ being 0, $1/2$, or $1$. To illustrate, when the density function $p(\cdot)$ is chosen to be that of a normal distribution and the parameter $\lambda$ is identified as the conditional variance $\sigma^{2}$, with $a=1$, the score-driven model reduces to the classical GARCH model.
The equation given by (2.1) represents a first-order score-driven model, as the relationship between the score and the parameter is lagged by only one period. This framework can be extended to higher orders, such as $p$-th order or even infinite order, as in janus2014long,
@@
It should be noted that our time index starts from zero; thus, the term “infinite” here refers to dependence on the entire history-from time 0 up to the present. Additionally, we omit the autoregressive term, as it can be absorbed into the coefficients $\phi_{i}$ through the moving average representation. This formulation appears similar to the ARCH($\infty$) model, but the key difference lies in the fact that the infinite-order lagged term $S\nabla$ in the score-driven model is a martingale difference, whereas the $y^{2}$ term in the ARCH($\infty$) model is not. In other words, rewriting the ARCH($\infty$) model in a form driven by martingale differences gives
@@
which includes an additional infinite autoregressive component. To address this, opschoor2019fractional proposed the following formulation:
@@
where $\phi_{i}^{(n)}=a_{n}\phi_{i}>0$. Let $a_{n}>0$ and $\uparrow 1$ as $n\to\infty$, $\|\phi\|=\sum_{i=1}^{\infty}\phi_{i}=1$, so that $\|\phi^{(n)}\|\uparrow 1$. This case is referred to as nearly unstable. Specifically, we set $\phi_{i}\sim K/i^{1+\alpha}$, exhibiting power-law decay, where $\alpha\in(1/2,1)$, a form referred to as “heavy-tailed” in jaisson2016rough. This condition is critical for the limit to generate roughness. The sigma-algebra is defined as $\mathcal{F}^{(n)}_{t}=\sigma(\lambda_{0}^{(n)},y_{i}^{(n)},\lambda_{i+1}^{(n)},i=0,1,\ldots,t)$, since $\lambda_{t+1}^{(n)}$ is determined when $\{y_{i}^{(n)}\}_{i\leq t}$ are known.
@@
where $\{W_{t}\}_{t\geq 0}$ is a standard Brownian motion. And according to cai2024scaling, in that case $1/2<\alpha<1$, when $1-a_{n}\sim n^{-\alpha}$ tends to zero and the coefficients $\phi_{i}\sim K/i^{1+\alpha}$ exhibit power-law decay, the following weak convergence holds:
@@
In fact, similar to Proposition 4.10 in el2018microstructural, the SDE (2.7) has the same solution as the SDE
@@
Since $\alpha-1\in(-1/2,0)$, the term $\frac{1}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}{\rm d}W_{s}$ corresponds to a fractional Brownian motion with Hurst parameter $H<1/2$, which means (2.8) falls into the rough case. It is worth mentioning that when $U(x)=\sqrt{x}$, this actually corresponds to the rough fractional CIR process described in jaisson2016rough.
@@ 3 Main results
The function $U:\mathbb{R}\to\mathbb{R}$ is locally bounded and converges to a finite limit $\gamma$ as $|x|\to\infty$.
@@
The martingale difference arrays $\{\Delta M_{k}^{(n)}\}_{k\leq k_{n},n\geq 1}$ satisfy: for every $p>2$,
$\sup_{n\geq 1}\max_{k\leq k_{n}}\mathbb{E}\left|\Delta M_{k}^{(n)}\right|^{p}<\infty.$
Remark 1.
Without loss of generality, we often set the terminal time $T=1$, so that $k_{n}=\lfloor nT\rfloor=n$ is a common and sufficient choice. This assumption controls all higher-order moments to ensure the validity of the following theorem for all $\alpha\in(1/2,1)$. In fact, if only a finite $p$-th moment exists for some $p>2$, the theorem still holds for $\alpha\in[\frac{p+2}{2p},1)$.
@@
In fact, we restrict our attention to the case of rough OU process in (2.8). This is because our proof strategy fundamentally differs from that of jaisson2016rough or wang2025rough: while they show that the Hawkes process (equivalent to the integrated intensity in probability) converges to a integrated rough CIR process, we directly establish the convergence of the parameter process itself to a rough process, rather than its integral. The difficulty in dealing with the process itself lies in verifying its tightness and we attempt to circumvent this by applying the continuous mapping theorem.
To this end, we regard the rough diffusion process as the image of a diffusion process under the generalized fractional operator (GFO), which is inspired by horvath2024functional. Specifically, for $\lambda\in(0,1)$ and $\alpha\in(-\lambda,1-\lambda)$, $\mathcal{G}^{\alpha}$ is the GFO associated to kernel
@@
$(\mathcal{G}^{\alpha}f)(t)=\left\{\begin{aligned} &\int_{0}^{t}(f(s)-f(0))\frac{{\rm d}}{{\rm d}t}g(t-s){\rm d}s,\quad\text{if}\ \alpha\geq 0,\\ &\frac{{\rm d}}{{\rm d}t}\int_{0}^{t}(f(s)-f(0))g(t-s){\rm d}s,\quad\text{if}\ \alpha<0.\end{aligned}\right$. (3.2)
When $g(x)=\frac{x^{\alpha}}{\Gamma(\alpha+1)}$, this reduces to the classical Riemann-Liouville operator. In this paper, we choose $g(x)=f^{\alpha,\kappa}(x)\in\mathcal{L}^{\alpha-1}$ and focus on the case $\alpha\in(1/2,1)$. Consequently, the corresponding GFO actually becomes
@@
$\Lambda_{t}:=\theta\int_{0}^{t}f^{\alpha,\kappa}(s){\rm d}s+\nu\gamma\int_{0}^{t}f^{\alpha,\kappa}(t-s){\rm d}W_{s}$ (3.4)
actually is the solution of the equation (3.1) we aim to converge to:
Proposition 2.
The process $\Lambda$ is the unique strong solution to the following rough SDE:
$\Lambda_{t}=\frac{1}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}\kappa(\theta-\Lambda_{s}){\rm d}s+\frac{\gamma\kappa\nu}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}{\rm d}W_{s}.$
Proof.
See A.1. ∎
Accordingly, in the subsequent proof, we take the clearer form of equation (3.4) as the target of convergence. Consider the partition $\mathcal{T}:=\{t_{i}=i/n,i=0,1,\ldots,n\}$. We then define
@@ 3.1 Converges to $f^{\alpha,\kappa}(x)$
jaisson2016rough prove that the mean of geometric sums converges in distribution to a random variable with density function $f^{\alpha,\kappa}(x)$. Our setting differs slightly because the summed terms follow a discrete distribution.
@@
where $\hat{\phi}$ denotes the Fourier transform of $\phi$, or equivalently, the characteristic function of the random variable $U_{1}$. Since $\phi_{i}\sim\frac{K}{i^{1+\alpha}}$, by the Karamata-Tauberian theorem (see cai2024scaling), we can characterize the behavior of $\hat{\phi}$ near the origin,
@@
See A.2. ∎
Utilizing the Fourier isometry, the estimates for the characteristic function $\hat{\phi}_{n}$ established in Lemma 3 help us analyze the $L^{2}$ distance $\|\rho_{n}-f^{\alpha,\kappa}\|_{2}$ in the Fourier domain. We are now equipped to obtain the stronger convergence result.
@@
$\|\rho_{n}-f^{\alpha,\kappa}\|_{2}:=\left(\int_{0}^{1}\left[\rho_{n}(x)-f^{\alpha,\kappa}(x)\right]^{2}{\rm d}x\right)^{1/2}\to 0,\ \text{as}\ n\to\infty.$ (3.9)
@@
See A.3. ∎